EVOLUTIONARY/GEOLOGICAL TIMELINE

Time (Myr ago) Event
4600  Formation of the approximately homogeneous solid Earth by planetesimal accretion
4300  Melting of the Earth due to radioactive and gravitational heating which leads to its differentiated interior structure as

well as outgassing of molecules such as water, methane, ammonia, hydrogen, nitrogen, and carbon dioxide

4300  Atmospheric water is photodissociated by ultraviolet light to give oxygen atoms which are incorporated into an ozone

layer and hydrogen molecules which escape into space

4300  First minerals (zircons) crystalize on the molten Earth
4000  Bombardment of the Earth by planetesimals stops
3800  The Earth’s crust solidifies—formation of the oldest rocks found on Earth
3800  Condensation of atmospheric water into oceans
3500–2800  Prokaryotic cell organisms develop
3500–2800  Beginning of photosynthesis by blue-green algae which releases oxygen molecules into the atmosphere and steadily works to

strengthen the ozone layer and change the Earth’s chemically reducing atmosphere into a chemically oxidizing one

2400  Rise in the concentration of oxygen molecules stops the deposition of uraninites (since they are soluble when combined

with oxygen) and starts the deposition of banded iron formations

2000  The Oklo natural fission reactors in Gabon go into operation and run for about 1 Myr
2000  The Magellanic Clouds tidally interact with the Milky Way thereby starting the Magellanic Stream (MNRAS 198, 710)
1600  The last reserves of reduced iron are used up by the increasing atmospheric oxygen—last banded iron formations
1500  Eukaryotic cell organisms develop
1500–600  Rise of multicellular organisms
1100  Formation of the supercontinent Rodinia
700  Break-up of the supercontinent Rodinia
580–545  Fossils of Ediacaran organisms are made
550  Formation of Gondwana
545  Cambrian explosion of hard-bodied organisms
528–526  Fossilization of the Chengjiang site
517–515  Fossilization of the Burgess Shale
500–450  Rise of the fish—first vertebrates
440  Ordovician mass extinction
430  Waxy coated algae begin to live on land
420  Millipedes have evolved—first land animals
375  The Appalachian mountains are formed via a plate tectonic collision between North America, Africa, and Europe
375  Appearance of primitive sharks
365  Devonian mass extinction
350–300  Rise of the amphibians
350  Primitive insects have evolved
350  Primitive ferns evolve—first plants with roots
300–200  Rise of the reptiles
300  Winged insects have evolved
280  Beetles and weevils have evolved
250  Permian period mass extinction
230  Roaches and termites have evolved
225  Modern ferns have evolved
225  Bees have evolved
210  Triassic mass extinction
200  Pangaea starts to break apart
200  Primitive crocodiles have evolved
200  Appearance of mammals
145  Archaeopteryx walks the Earth
136  Primitive kangaroos have evolved
100  Primitive cranes have evolved
90  Modern sharks have evolved
65  The Chicxulub impact occurs
65  K-T Boundary—extinction of the dinosaurs and beginning of the reign of mammals
60  Rats, mice, and squirrels have evolved
60  Herons and storks have evolved
55  Rabbits and hares have evolved
50  Primitive monkeys have evolved
28  Koalas have evolved
20  Parrots and pigeons have evolved
20–12  The chimpanzee and hominid lines evolve
10–4  Ramapithecus exist
4  Development of hominid bipedalism
4–1  Australopithecus exist
3.5  The Australopithecus Lucy walks the Earth
2  Widespread use of stone tools
2–0.01  Most recent ice age
1.6–0.2  Homo erectus exist
1–0.5  Homo erectus tames fire
0.3  Geminga supernova explosion at a distance of roughly 60 pc—roughly as bright as the Moon
0.2–0.03  Homo sapiens neanderthalensis exist
0.05–0  Homo sapiens sapiens exist
0.04–0.012  Homo sapiens sapiens enter Australia from southeastern Asia and North America from northeastern Asia
0.025–0.01  Most recent glaciation—an ice sheet covers much of the northern United States
0.02  Homo sapiens sapiens paint the Altamira cave and the Chauvet cave
0.017  Homo sapiens sapiens paint the Lascaux cave
0.012  Homo sapiens sapiens have domesticated dogs in Kirkuk, Iraq

0.01  First permanent homo sapiens sapiens settlements
0.01  Homo sapiens sapiens learn to use fire to cast copper and harden pottery
0.006  Writing is developed in Sumeria
0.0046  Oldest known Bristlecone Pine tree starts to grow

Time (Myr ago) Event
5000–1500  Archeozoic (Archean) era
1500–545  Proterozoic era

Paleozoic era “Ancient life”

545–505  Cambrian period
505–438  Ordovician period
438–410  Silurian period
410–355  Devonian period
355–290  Carboniferous (Mississippian/Pennsylvanian) period
290–250  Permian period

Mesozoic era “Middle Life”

250–205  Triassic period
205–135  Jurassic period
135–65  Cretaceous period

Cenozoic era “Recent Life”

Tertiary (Paleogene/Neocene) period

65–55  Paleocene epoch
55–38  Eocene epoch
38–26  Oligocene epoch
26–6  Miocene epoch
6–1.8  Pliocene epoch

Quaternary period

1.8–0.01  Pleistocene epoch
0.50–0.25  Lower Paleolithic
0.25–0.06  Middle Paleolithic
0.06–0.01  Upper Paleolithic
0.01–0  Holocene epoch